2. Given z1 = 2∠45o


, ; z2 = 3∠120o and z3 = 4∠180o


. Determine the following and leave your


answers in rectangular form:


(i)


(z1)2 +z2


z2 +z3


(5)


(ii)


z1


z2z3


1
Expert's answer
2022-06-03T11:25:02-0400

(i)

"(z_1)^2=(2)^2\angle(2\cdot45\degree)=4i"

"z_2+z_3=3(-\dfrac{1}{2}+\dfrac{\sqrt{3}}{2}i)-4=-\dfrac{11}{2}-\dfrac{3\sqrt{3}}{2}i)"

"\dfrac{z_2}{z_2+z_3}=\dfrac{-3+3\sqrt{3}i}{-11-3\sqrt{3}i}"

"=\dfrac{(-3+3\sqrt{3}i)(-11-3\sqrt{3}i)}{121+27}"


"=\dfrac{33+9\sqrt{3}i-33\sqrt{3}i+27}{148}"


"=\dfrac{15}{37}-\dfrac{6\sqrt{3}}{37}i"


"(z_1)^2+\dfrac{z_2}{z_2+z_3}=4i+\dfrac{15}{37}-\dfrac{6\sqrt{3}}{37}i"

"=\dfrac{15}{37}+\dfrac{148-6\sqrt{3}}{37}i"

(ii)


"z_2z_3=3(4)\angle(120\degree+180\degree)=6-6\sqrt{3}i"

"\dfrac{z_1}{z_2z_3}=\dfrac{\sqrt{2}+\sqrt{2}i}{6-6\sqrt{3}i}"

"=\dfrac{(\sqrt{2}+\sqrt{2}i)(6+6\sqrt{3}i)}{36+108}"

"=\dfrac{6\sqrt{2}+6\sqrt{6}i+6\sqrt{2}i-6\sqrt{6}}{144}"

"=-\dfrac{\sqrt{6}-\sqrt{2}}{24}+\dfrac{\sqrt{6}+\sqrt{2}}{24}i"